a 12 lit. solution in 33.33 %acid . how much water must be added to get the sol. having 20% acid
Answer by mananth(16946)

--- percent ---------------- Amount
Acid solution 33.33 ---------------- 12 liters
Water 0 ---------------- x liters
Mixture 10 ---------------- 12 + x liters

33.33 * 12 + 0 x = 10 ( 12 + x )
399.96 + 0 x = 120 + 10 x
0 x -10 x = -399.96 + 120
-10 x = -279.96
/ -10
x = 28 liters Water

m.ananth@hotmail.ca